OpenStudy (anonymous):

Can someone please help me with this problem i have been stuck on it forever Raise the quantity in parentheses to the indicated exponent, and simplify the resulting expression. Express answers with positive exponents. (-39x^-2y/156x^3y^-3)^3

OpenStudy (amistre64):

cube it all

OpenStudy (amistre64):

(-39x^-2y/156x^3y^-3)^3 -39^3 x^-2^3 y^3 / 156^3 x^3^3 y^-3^3 if an exponent is cubed; just mulitply them together -39^3 x^-6 y^3 / 156^3 x^9 y^-9

OpenStudy (amistre64):

if you got a negative exponent, put it to the other side of the fraction bar to positify it

OpenStudy (amistre64):

and then simplfy as needed

OpenStudy (asnaseer):

\[(\frac{-39x^{-2}y}{156x^{3}y^{-3}})^{3}\]remember that a^-n means 1/a^n and it to 1st remove all negative powers:\[(\frac{-39y^{4}}{156x^{5}})^{3}\]simplify the constants by dividing by the common factor (39):\[(\frac{-y^{4}}{4x^{5}})^{3}\]now just cube the result:\[\frac{-y^{12}}{64x^{15}}\]